OpenStudy (sleepyjess):

Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x. f(x) = x3 + 4 and g(x) = Cube root of quantity x minus four.

OpenStudy (sleepyjess):

@Compassionate @johnweldon1993

OpenStudy (johnweldon1993):

Well we would just plug them into each other \[\large f(g(x)) = (\sqrt[3]{x-4})^3 + 4\] \[\large f(g(x)) = (x - 4) + 4\] \[\large f(g(x)) = x\] and for the other \[\large g(f(x)) = \sqrt[3]{(x^3 + 4) - 4}\] \[\large g(f(x)) = \sqrt[3]{(x^3)}\] \[\large g(f(x)) = x^{3/3} = x\] \[\large \color \red{\checkmark}\]

OpenStudy (sleepyjess):

Thank you so much! I knew what I had to do (plug them into each other) but just wasn't sure how to do it

OpenStudy (sleepyjess):

@johnweldon1993 could you help me on 1 more?

OpenStudy (johnweldon1993):

Sure, apologies in advance for any delays as I am doing work at the same time lol

OpenStudy (sleepyjess):

Its fine :-)

OpenStudy (sleepyjess):

f(x) = 3x + 6, g(x) = 2x^2 Find (fg)(x). 2x^2 + 3x + 6 6x + 12 6x^2 + 12x 6x3 + 12x^2

OpenStudy (johnweldon1993):

Okay, so noice how the grouping is different here \[\large (fg)(x)\] so this means we need to multiply our 2 functions... \[\large f = 3x + 6\] \[\large g(x) = 2x^2\] so \[\large (fg) = (3x + 6) \times (2x^2)\]

OpenStudy (sleepyjess):

6x^3+12x^2 ?

OpenStudy (johnweldon1993):

Correct indeed!